OpenStudy (anonymous):

Part 1: Solve each of the quadratic equations below and describe what the solution(s) represent to the graph of each. Show your work to receive full credit. x2 - 36 = 0 and x2 = 8x - 12 Part 2: Using complete sentences, compare and contrast the graphs of y = x2 − 36 and y + x2 = 8x − 12.

OpenStudy (anonymous):

x^2 - 36 = 0 I can see immediately that x^2 - 36 = (x-6)(x+6) so then x can be 6 or -6.

OpenStudy (zale101):

I'll do the first one for you because it looks easy. All you have to do is take 36 to the other side of the equals. x^2-36=0 with difference of squares (x+6)(x-6) x^2 - 36 = 0-> x^2 = 36 x = +/- 6 (so it can be positive 6 or negative 6)

OpenStudy (anonymous):

For the second equation, start by making it a quadratic equation: x^2 - 8x + 12 = 0 Then you want to factor it by finding a, b such that a + b = -8 and ab = 12

OpenStudy (anonymous):

Thank you.